OpenStudy (anonymous):

A piece of wire 5 inches long is to be cut into two pieces. One peice is x inches long and is to be bent into the shape of a square. The other piece is to be bent into the shape of a circle. Find an expression for the total adrea made up by the square and the circle as a function of x.

OpenStudy (campbell_st):

for the square Perimeter = x side length = x/4 then area A = x^2/16 The circle if x inches in used there is 5 - x inches remaining so the circumference = 5 -x find the radius \[5 - x = 2\pi r\] \[r = \frac{5 - x}{\pi}\] area of the circle \[A = \pi \times(\frac{5 -x}{\pi})^2\] \[A = \frac{1}{\pi} (5 - x)^2\] or \[A = \frac{1}{\pi} 25 - 10x + x^2\] then the total area is \[f(x) = \frac{x^2}{16} + \frac{1}{\pi}( 25 - 10x + x^2)\] giving the final function as \[f(x) = (\frac{1}{16} + \frac{1}{\pi})x^2 -10x + 25\]

OpenStudy (campbell_st):

you many want to simplify the coefficient of x^2

OpenStudy (anonymous):

what happened to the 2 when you divided 2π to find r?

OpenStudy (campbell_st):

oops missed a 2 the radius should be \[r = \frac{5 -x}{2\pi}\] so area of the circle is \[A = \pi (\frac{(5 -x)}{2\pi})^2\] \[A = \frac{(5 - x)^2}{4\pi}\] so the total area is \[f(x) = \frac{x^2}{16} + \frac{(5 - x)^2}{2\pi}\] I'll let you simplify it... sorry for the error in the 1st posting

OpenStudy (campbell_st):

dang... total area \[f(x) = \frac{x^2}{16} + \frac{(5 -x)^2}{4\pi}\] which is finally correct... sorry for the typos

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

close this ..